OpenStudy (anonymous):

The surface areas of two similar solids are 450mm squared and 578mm squared. If the volume of the smaller solid is 511mm cubed find the volume of the larger solid. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

if the sides of Solid A is K times bigger than Solid B Then the area will be k^2 times more and the volume will be k^3 times more

OpenStudy (dan815):

578=k^2*450, solve for k Volume of bigger solid=k^3*Volume of smaller solid =k^3*511

OpenStudy (anonymous):

Brooks built two similar wooden chests in the shape of rectangular prisms. She used 5 pints of paint to cover the outside of the larger chest fully and 2 pints to cover the outside of the smaller chest. How much more volume does the larger chest contain?

OpenStudy (anonymous):

thank you! can you help me on the previous one? @dan815 ^^^^^^

OpenStudy (dan815):

Rectangular Prisms, you are given both of their surface area relationship, indirectly, and they want you to find the ratio between their volumes. It is a similar type of question as before, try to work on it, on your own.